OpenStudy (anonymous):

how do you find the equation for (2,4) and (3,1)

OpenStudy (aroub):

first u have to find the slope (do u know how?)

OpenStudy (anonymous):

yes i believe its 5

OpenStudy (aroub):

ok let me check, 1-4/3-2 = -3/1 =-3

OpenStudy (aroub):

how do u find the slope??

OpenStudy (aroub):

or how did you*** sorryy

OpenStudy (anonymous):

i did -1 -4 / 3 -2 = -5/1 x -1 = 5

OpenStudy (aroub):

why did u do the 1 negative??

OpenStudy (anonymous):

the one in (3, 1) is actually negative sorry

OpenStudy (aroub):

oh okay lol! so the slope is -5

OpenStudy (anonymous):

oh kat so what do i do after that?

OpenStudy (aroub):

okay u pick of any of the points u have ( ill pick (2,4) ) to find the y-intercept for the equation ill show u how! :)

OpenStudy (anonymous):

yay thank you soo much

OpenStudy (aroub):

y=mx+b u want to find the B ok? ok ^ ^ slope y-intercept <--- i guess u know this right? (lol ill write the y=mx+b alot so just dont get surprised) y=mx+b , m(slope)=-5 , b=(2,4) (x,y) ok now u have to substitute the 2 by x and the 4 by y to find the y-intercept y=mx+b ^ ^ ^ 4=-5(2)+b , ( now u just have to find B) 4=-10+B 10+4=B 14=B , (OK THE Y-INTERCEPT IS 14) so the equation u have is y=-5x+14 did u understand anything? if yes im so gglllaaddd!!! / if no just tell me what u didnt understand!! :)

OpenStudy (anonymous):

i get it thank uuu sooo much :))

OpenStudy (aroub):

no problem chlojo7!!! :D :D